# On the existence of weighted-cscK metrics

### v2

@@ 3.C. Gradient estimate of $F-\log\mathrm{v}+f_{*}$
where $C$ is a constant depending on $\omega$, $[\theta]$, $\mathrm{v}$, $\mathrm{w}$, $\|\phi\|_{C^{0}}$, $\|F+f_{*}\|_{C^{0}}$, $\|\mathrm{tr}\omega_{\phi}\|_{L^{p}}$ for some $p$.
@@
Since $f_{*}$ is a $(\frac{1-t}{t})\theta$-psh function , $\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by class $[\theta]$, which is bounded. And by Theorem 3.2, $|W|_{C^{0}}$ is bounded. With the preliminaries in mind and under an orthonormal frame of $g_{\phi}$, using Bochner formula (see the proof of [CC21a, Proposition 4.1]) implies
@@
$=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F^{\prime}+f_{*})-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})+\left(\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}-\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$ (3.15)
@@
$2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)-C-C\mathrm{tr}_{\phi}\omega-C\mathrm{tr}\omega_{\phi}\leq\Delta_{\phi}W\leq 2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+C+C\mathrm{tr}_{\phi}\omega+C\mathrm{tr}\omega_{\phi},$
where we use (3.4) and (2.16) for the second equality, (2.18) for the last inequality.
@@
$+e^{-\log\mathrm{v}(m_{\phi})}\left(\left(\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(W)-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(\phi)}J\xi^{\alpha}(f_{*})-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})\right.\right.$
@@
For any $20<\varepsilon<$ 2, by Hölder inequality, we have
@@ 4. Proof of Continuity Method
$\int_{X}\mathrm{tr}_{\mathrm{v},\omega}(\theta)\frac{\omega^{n}}{n!}=\int_{X}\mathrm{v}(m_{\omega})\theta\wedge\frac{\omega^{n-1}}{(n-1)!}+\int_{X}\langle d\mathrm{v}(m_{\omega}),m_{\theta}\rangle\frac{\omega^{n}}{n!}=\underline{\theta}\cdot\int_{X}\mathrm{w}(m_{\omega})\frac{\omega^{n}}{n!}.$
@@ 4.A. Openness
In this subsection, we show the openness of the continuity path (1.3), following the standard implicit function theory ([Che18], [Has19]). Consider
@@
~$\widetilde{{\mathcal{F}}}:{\mathcal{H}}^{4,\alpha}_{K,0}\times[0,1]$ $\rightarrow C^{\alpha}(X)_{K,0}$
@@
Note that $\int_{X}{\mathcal{F}}(\phi,t)\omega_{\phi}^{n}=0$. Therefore if ~$\widetilde{{\mathcal{F}}}(\phi,t)=0$, ${\mathcal{F}}(\phi,t)$=0. Thus it suffices to apply implicit function theorem on the map $\widetilde{{\mathcal{F}}}$ near $(\phi_{0},t_{0})$. Denote
(ϕ):=assignitalic-$\mathcal{R}(\phi):=$ $\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\rm ext}(m_{\phi}),\text{ and }\mathcal{T}(\phi):=\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}.$
@@
where $\mathcal{D}_{\phi}u:=\sqrt{2}(\nabla^{\phi}\mathrm{d}u)^{-}$ is the $J$-anti-invariant part of the tensor $(\nabla^{\phi}\mathrm{d}u)$, with $\nabla^{\phi}$ the $g_{\phi}$-Levi-Civita connection, and $\mathcal{D}^{*}_{\phi}$ is the formal adjoint of $\mathcal{D}_{\phi}$.
@@
By Lemma 5.4, we have that the operator ${\mathcal{L}}_{(\phi_{0},t_{0})}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi}$-self-adjoint $K$-invariant linear operator satisfying
$({\mathcal{L}}_{(\phi_{0},t_{0})}u_{1},u_{2})_{\mathrm{w},\phi}=$ $\int_{X}{\mathcal{L}}_{(\phi_{0},t_{0})}u_{1}\cdot u_{2}\mathrm{w}(m_{\phi_{0}})\omega_{\phi_{0}}^{n}$
@@
$=$ $(u_{1},{\mathcal{L}}_{(\phi_{0},t_{0})}u_{2})_{\mathrm{w},\phi}.$
When $t_{0}>0$, following the argument in [Che18], we divide the remaining argument into three steps:
@@
The proof follows from [Che18, Lemma 4.3] by using Schauder estimates, Poincaré inequality.
@@
Similar to [Che18, Lemma 4.4], the proof is given by a standard Fredholm theorem for elliptic operators.
@@
$\qquad-\int_{X}\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})-r\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)\omega_{\phi_{0}}^{n}$
@@
${\mathcal{L}}_{0,\phi_{0}}(u):=(D_{\phi_{0}}{\mathcal{F}}_{0})(u)=-\frac{(\mathcal{D}^{*}_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\mathcal{D}_{\phi_{0}})u}{\mathrm{w}(m_{\phi_{0}})}+rF_{\phi_{0}}(u),$
@@
which is a $(\cdot,\cdot)_{\mathrm{w},\phi}$-self-adjoint $K$-invariant linear operator with a trivial kernal (see Lemma 5.4).
@@
By the detailed computation of the linearized operators in Appendix 5.B, we find that $F_{\phi_{0}}$ has a negative first eigenvalue. Thus, for $r$ sufficiently large, the first eigenvalue $\lambda_{1,r}$ of ${\mathcal{L}}_{0,\phi_{0}}$ is always negative.
@@
At the end, the implicit function theorem argument in [Has19] implies that there exists a solution $\phi\in{\mathcal{H}}_{K}$ of ${\mathcal{F}}_{0}(\phi)=0$ for $r$ large enough.
Combining the two cases of $t_{0}=0$ and $10<t_{0}<$ 1, we obtain the following openness property.
@@ 4.B. Closedness
    The weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ is bounded from below. Furthermore, ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$ is coercive for $10<t<$ 1.

### v3

@@ 3.C. Gradient estimate of $F-\log\mathrm{v}+f_{*}$
where $C$ is a constant depending on $\omega$, $\theta]$, $\mathrm{v}$, $\mathrm{w}$, $\|\phi\|_{C^{0}}$, $\|F+f_{*}\|_{C^{0}}$, $\|\mathrm{tr}\omega_{\phi}\|_{L^{p}}$ for some $p$.
@@
Since $f_{*}$ is a $(\frac{1-t}{t})\theta$-psh function , $\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by class $\theta]$, which is bounded. And by Theorem 3.2, $|W|_{C^{0}}$ is bounded. With the preliminaries in mind and under an orthonormal frame of $g_{\phi}$, using Bochner formula (see the proof of [CC21a, Proposition 4.1]) implies
@@
$=$ $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(F^{\prime}+f_{*})-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})+\left(\frac{\mathrm{v}_{,\alpha}\mathrm{v}_{,\beta}}{\mathrm{v}^{2}}-\frac{\mathrm{v}_{,\alpha\beta}}{\mathrm{v}}\right)\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$
@@
which together with (2.19) implies
$2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)-C-C\mathrm{tr}_{\phi}\omega-C\mathrm{tr}\omega_{\phi}\leq\Delta_{\phi}W\leq 2\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(W)+C+C\mathrm{tr}_{\phi}\omega+C\mathrm{tr}\omega_{\phi}$ (3.17)
@@
[…]
@@
For any $0<\varepsilon<2$, by Hölder inequality, we have
@@ 4. Proof of Continuity Method
and $\underline{\theta}$ is a contant determined by the equality:
$\int_{X}\mathrm{tr}_{\mathrm{v},\omega}(\theta)\frac{\omega^{n}}{n!}=\int_{X}\mathrm{v}(m_{\omega})\theta\wedge\frac{\omega^{n-1}}{(n-1)!}+\int_{X}\langle\mathrm{d}\mathrm{v}(m_{\omega}),m_{\theta}\rangle\frac{\omega^{n}}{n!}=\underline{\theta}\cdot\int_{X}\mathrm{w}(m_{\omega})\frac{\omega^{n}}{n!}.$ (4.2)
In particular, $\underline{\theta}$ only depends on $\omega]$ ([Lah19, Lemma 3]).
@@ 4.A. Openness
In this subsection, we show the openness of the continuity path (1.3), following the standard implicit function theory ([Ch18], [Has19]). Firstly, note that as explained in the proof of Proposition 4.2 (1), we can choose $K$-invariant $\theta\in[\omega]$ such that $\phi=0$ is the solution of continuity path at $t=0$.
Consider
@@
$\widetilde{{\mathcal{F}}}:{\mathcal{H}}^{4,\alpha}_{K,0}\times[0,1]$ $\rightarrow C^{\alpha}(X)_{K,0}$
@@
Note that $\int_{X}{\mathcal{F}}(\phi,t)\mathrm{w}(m_{\phi})\omega_{\phi}^{n}=0$ by the definition of the constant $\underline{\theta}$ and extremal function $\ell_{\mathrm{ext}}$. Therefore if $\widetilde{{\mathcal{F}}}(\phi,t)=0$, ${\mathcal{F}}(\phi,t)$=0. This is because $\widetilde{{\mathcal{F}}}(\phi,t)=0$ implies ${\mathcal{F}}(\phi,t)\equiv C$, then by integrating with respect to the weighted measure $\mathrm{w}(m_{\phi})\omega_{\phi}^{n}$, we obtain that $C$ must be zero. Thus it suffices to apply implicit function theorem on the map $\widetilde{{\mathcal{F}}}$ near $(\phi_{0},t_{0})$. Denote
$\mathcal{R}(\phi):=$ $\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\rm ext}(m_{\phi}),\text{ and }\mathcal{T}(\phi):=\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}.$
@@
where $\mathcal{D}_{\phi}u:=\sqrt{2}(\nabla^{\phi}\mathrm{d}u)^{-}$ is the $J$-anti-invariant part of the tensor $(\nabla^{\phi}\mathrm{d}u)$, with $\nabla^{\phi}$ the $g_{\phi}$-Levi-Civita connection, and $\mathcal{D}^{*}_{\phi}$ is the formal adjoint of $\mathcal{D}_{\phi}$ with respect to $(\cdot,\cdot)_{\phi}$, where
$(f,h)_{\phi}=\int_{X}f\;h\;\omega_{\phi}^{n},$
for any $f,h\in C^{4,\alpha}(X)_{K,0}$.
@@
By Lemma 5.4, we have that the operator ${\mathcal{L}}_{(\phi_{0},t_{0})}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi_{0}}$-self-adjoint $K$-invariant linear operator satisfying
$({\mathcal{L}}_{(\phi_{0},t_{0})}u_{1},u_{2})_{\mathrm{w},\phi_{0}}=$ $\int_{X}{\mathcal{L}}_{(\phi_{0},t_{0})}u_{1}\cdot u_{2}\mathrm{w}(m_{\phi_{0}})\omega_{\phi_{0}}^{n}$
@@
$=$ $(u_{1},{\mathcal{L}}_{(\phi_{0},t_{0})}u_{2})_{\mathrm{w},\phi_{0}}.$
When $t_{0}>0$, following the argument in [Ch18], we divide the remaining argument into three steps:
@@
The proof follows from [Ch18, Lemma 4.3] by using Schauder estimates, Poincaré inequality.
@@
Similar to [Ch18, Lemma 4.4], the proof is given by a standard Fredholm theorem for elliptic operators.
@@
$\hskip 18.49988pt-\int_{X}\frac{S_{\mathrm{v}}(\phi)}{\mathrm{w}(m_{\phi})}-\ell_{\mathrm{ext}}(m_{\phi})-R\left(\frac{\mathrm{tr}_{\mathrm{v},\phi}(\theta)}{\mathrm{w}(m_{\phi})}-\underline{\theta}\right)\omega_{\phi_{0}}^{n}$
@@
${\mathcal{L}}_{0,\phi_{0}}(u):=(D_{\phi_{0}}{\mathcal{F}}_{0})(u)=-\frac{(\mathcal{D}^{*}_{\phi_{0}}\mathrm{v}(m_{\phi_{0}})\mathcal{D}_{\phi_{0}})u}{\mathrm{w}(m_{\phi_{0}})}+RF_{\phi_{0}}(u),$ (4.5)
@@
which is a $(\cdot,\cdot)_{\mathrm{w},\phi_{0}}$-self-adjoint $K$-invariant linear operator with a trivial kernal (see Lemma 5.4).
@@
By the detailed computation of the linearized operators in Appendix 5.B, we find that $F_{\phi_{0}}$ has a negative first eigenvalue. Thus, the first eigenvalue $\lambda_{1,R}$ of ${\mathcal{L}}_{0,\phi_{0}}$ is always negative for $R>0$. Indeed, let $\phi_{1,R}$ be the first eigenfunction of ${\mathcal{L}}_{0,\phi_{0}}$, one has
$\lambda_{1,R}=$ $\frac{({\mathcal{L}}_{0,\phi_{0}}(\phi_{1,R}),\phi_{1,R})_{\mathrm{w},\phi_{0}}}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}$
$=$ $\frac{1}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}\left(-\int_{X}\phi_{1,R}({\mathcal{D}}_{\phi_{0}}^{*}\mathrm{v}(m_{\phi_{0}}){\mathcal{D}}_{\phi_{0}})\phi_{1,R}\omega_{\phi_{0}}^{n}+R(F_{\phi_{0}}(\phi_{1,R}),\phi_{1,R})_{\mathrm{w},\phi_{0}}\right)$
$=$ $\frac{-1}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}\int_{X}\mathrm{v}(m_{\phi_{0}})|{\mathcal{D}}_{\phi_{0}}\phi_{1,R}|^{2}_{\phi_{0}}\omega_{\phi_{0}}^{n}+R\frac{(F_{\phi_{0}}(\phi_{1,R}),\phi_{1,R})_{\mathrm{w},\phi_{0}}}{(\phi_{1,R},\phi_{1,R})_{\mathrm{w},\phi_{0}}}$
$\leq$ $R\lambda_{1}(F_{\phi_{0}})\leq-C\cdot R<0.$ (4.6)
@@
At the end, the implicit function theorem argument in [Has19] implies that there exists a solution $\phi\in{\mathcal{H}}_{K}$ of ${\mathcal{F}}_{0}(\phi)=0$ for $R$ large enough. We sketch the main steps of Hashimoto’s argument as follows.
  1. (i)
    As stated in (4.8), we can find a solution $\omega$ for the continuity path at $t=0$. Perturb $\omega$ to get
    $\omega_{m}:=\omega+{\rm dd^{c}}(R^{-1}\phi_{1}+\cdots+R^{-m}\phi_{m})$
    for some $\phi_{1},\ldots,\phi_{m}\in C^{\infty}(X)_{K}$, satisfying ${\mathcal{R}}(\omega_{m})-R{\mathcal{T}}(\omega_{m})=const+R^{-m}f_{m,R}$ for a function $f_{m,R}\in C^{\infty}(X)_{K}$ with $(f_{m,R},1)_{\mathrm{w},\phi_{0}}=0$ which is bounded in $C^{\infty}(X)_{K}$ for all sufficiently large $R$. The proof follows from [Has19, Lemma 3].
    Then by solving a weighted Green’s function $\frac{\mathrm{d}^{*}(\mathrm{v}(m_{\omega_{m}})\mathrm{d}G_{m,R})}{\mathrm{w}(m_{\omega_{m}})}=f_{m,R}$, we can find that $\omega_{m}$ satisfies
    ${\mathcal{R}}(\omega_{m})-R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m}}\theta_{m}}{\mathrm{w}(m_{\omega_{m}})}=const$ (4.7)
    where $\theta_{m}:=\theta+R^{-m-1}{\rm dd^{c}}G_{m,R}$.
  2. (ii)
    Consider the Banach spaces $B_{1}=\Omega^{1,1}(X)_{K}\times L^{2}_{p+4}(X)_{K,0}$ and $B_{2}=\Omega^{1,1}(X)_{K}\times L^{2}_{p}(X)_{K,0}$, and an open set $U:=\{(\chi,\phi)\in B_{1}~|~\omega_{m,\phi}:=\omega_{m}+{\rm dd^{c}}\phi>0\}$, where $\Omega^{1,1}(X)_{K}$ is the $L^{2}_{p}(X)$ Sobolev completion of the space of $K$-invariant real $(1,1)$-forms and $L^{2}_{p}(X)_{K,0}$ is the Sobolev completion of $C^{\infty}(X)_{K}$ with $(\phi,1)_{\mathrm{w},\phi_{0}}=0$. We can define a map $T:U\rightarrow B_{2}$ by
    $T(\chi,\phi):=\left(\theta_{m}+\chi,{\mathcal{R}}(\omega_{m,\phi})-R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m,\phi}}(\theta_{m}+\chi)}{\mathrm{w}(m_{\omega_{m,\phi}})}\right)$
    Using (4.7), the linearization of $T$ at 0 is given by
    $(D_{0}T)(\chi,u)=\left(\begin{array}[]{cc}1&0\\ -R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m}}}{\mathrm{w}(m_{\omega_{m}})}&{\mathcal{L}}_{0,\omega_{m}}\end{array}\right)\left(\begin{array}[]{c}\chi\\ u\end{array}\right),$
    where ${\mathcal{L}}_{0,\omega_{m}}$ is defined as in (4.5), where we substitute $\theta$ by $\theta_{m}$ and substitute $\omega_{\phi_{0}}$ by $\omega_{m}$.
    By Appendix 5.B, we have ${\mathcal{L}}_{0,\omega_{m}}$ is a $(\cdot,\cdot)_{\mathrm{w},\omega_{m}}$-self-adjoint linear operator with trivial kernel. Then by Fredholm theory, we have that $D_{0}T:B_{1}\rightarrow B_{2}$ is an isomorphism with the inverse
    $P=\left(\begin{array}[]{cc}1&0\\ \left({\mathcal{L}}_{0,\omega_{m}}\right)^{-1}\left(\frac{-R\mathrm{tr}_{\mathrm{v},\omega_{m}}}{\mathrm{w}(m_{\omega_{m}})}\right)&\left({\mathcal{L}}_{0,\omega_{m}}\right)^{-1}\end{array}\right).$
    By the computation of (4.6) and Appendix 5.B (see also [Has19, lemma 4]), there is a constant $C=C(\omega,\theta)>0$ such that $\lambda_{1,m}<-CR$ for all sufficiently large $R$, where $\lambda_{1,m}<0$ is the largest non-zero eigenvalue of ${\mathcal{L}}_{0,\omega_{m}}$. By the above definition of $P$, one has $\|P\|_{\rm op}\leq CR^{2}$ (see [Has19, Below of Lemma 4]).
  3. (iii)
    By the definition of $T$, for $l\in{\mathbb{N}}$ and $l\geq 5$, for any $R$ sufficiently large, the higher order part $T-D_{0}T$ is Lipschitz with Lipschitz constant $1/(2\|P\|_{\rm op})$ in a ball centered at 0 of radius $\delta^{\prime}=R^{-l}$.
  4. (iv)
    We choose $\delta=\frac{\delta^{\prime}}{2\|P\|_{\rm op}}>\frac{1}{2C}R^{-l-2}$ for latter use.
  5. (v)
    Note that $T(0,0)=(\theta_{m},-R\underline{\theta})$. Since $\theta_{m}-\theta=O(R^{-m-1})$, we obtain
    $\|T(0,0)-(\theta,-R\underline{\theta})\|_{L^{2}_{p}}\leq CR^{-m-1}\leq\frac{1}{2C}R^{-l-2}<\delta,$
    for all large enough $R>0$, by taking $m$ to be sufficiently large. Choose $p>0$ to be sufficiently large such that $C^{4}(X)\hookrightarrow L^{2}_{p}(X)$. By the argument above, the assumptions of the quantitative inverse function theorem (see [Has19, Theorem 2]) are satisfied. By applying this theorem, there exists $(\chi,\phi)\in U$ such that $\theta_{m}+\chi=\theta$ and
    ${\mathcal{R}}(\omega_{m,\phi})-R\frac{\mathrm{tr}_{\mathrm{v},\omega_{m,\phi}}(\theta)}{\mathrm{w}(m_{\omega_{m,\phi}})}=R\underline{\theta}$
    for all large enough $R>0$. The standard elliptic PDE theory shows that $\phi$ is smooth.
Combining the two cases of $t_{0}=0$ and $0<t_{0}<1$, we obtain the following openness property.
@@ 1. $(1)$
    The weighted Mabuchi functional ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}}}$ is bounded from below. Furthermore, ${\mathbb{M}}_{\mathrm{v},\mathrm{w}\cdot\ell_{\mathrm{ext}},t}$ is coercive for $0<t<1$.